accelerates uniformly from a velocity of 1 m s-1 to a velocity of 5 m s-1 in 0.5 minutes . Calculate the displacement of the transporter .



1
Expert's answer
2021-02-18T18:55:29-0500
"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\cdot(1.0\ \dfrac{m}{s}+5.0\ \dfrac{m}{s})\cdot30\ s=90\ m."